# Unveiling Group-Specific Distributed Concept Drift: A Fairness Imperative in Federated Learning

### v3

@@ VI-A Results Across Multiple Datasets and Scenarios
VI-B Unveiling the Influence of $\alpha$ on Loss Over Time
@@
Figure 3: Average clients’ loss ($\ell$) and group-specific ($S=0$) loss ($\ell_{S=0}$) results for the MNIST-GDrift dataset considering $\alpha=0.1$ and $\alpha=0.5$.
@@
Figure 4: Clustering of FedDrift, FairFedDrift and Oracle on the MNIST-GDrift dataset, considering $\alpha=0.1$. Each cell represents the model ID at each client and timestep, while background colors denote the ground-truth concept.
We delve into the influence of group-specific distributed concept drift on the loss, emphasizing scenarios where the overall loss remains relatively stable while the loss of the unprivileged group experiences a noticeable decrease.
Figure 3 illustrates the loss ($\ell$) and group-specific ($S=0$) loss ($\ell_{S=0}$) results obtained from running the FedAvg algorithm on the MNIST-GDrift dataset considering $\alpha=0.1$ and $\alpha=0.5$. We showcase the results for the FedAvg since both FedDrift and FairFedDrift are built on top of this base algorithm.
@@
The results for $\alpha=0.1$ reveal that the loss of the unprivileged group varies significantly over time while the overall loss remains relatively stable. In such scenarios, FedDrift can struggle to detect significant loss changes, making the tuning of $\delta$ challenging since smaller values can result in too many clusters, while larger values may fail to detect group-specific concept drift. This highlights the limitations of FedDrift in addressing subtle, group-specific changes, as demonstrated before.
Conversely, when $\alpha=0.5$, both the overall loss and the loss of the unprivileged group exhibit noticeable variations. In these cases, FedDrift can successfully detect loss changes and create clusters accordingly, showcasing its capability in scenarios where the unprivileged group is well-represented and its losses are integral indicators of concept drift.
@@
Figure 4 presents the clustering outcomes achieved by FedDrift, FairFedDrift and Oracle on the MNIST-GDrift dataset considering $\alpha=0.1$. Although both FedDrift and FairFedDrift can detect concept drift in the third timestep, it can be observed in the fourth timestep that clients in the green concept in FedDrift prefer to associate themselves with the original model (‘0’), while FairFedDrift’s clients in the green concept associate themselves with models created in the previous timestep. While FairFedDrift’s approach to associating clients according to the sum of losses (line 8 in Algorithm 1) is crucial in this scenario, FedDrift lacks this mechanism, resulting in clients preferring the original model. In addition, it can be observed that FairFedDrift successfully merges models corresponding to the same concept (e.g. ‘1-3-4-5’). The visual analysis underscores FairFedDrift’s ability to achieve more accurate and adaptive clustering, thereby offering a robust solution this problem.
@@
Finally, FairFedDrift offers easier tuning in situations of group-specific concept drift compared to FedDrift. The loss of the unprivileged group undergoes significant variations over time, while the overall loss remains relatively stable. This makes tuning FedDrift more challenging, as it may struggle to detect significant loss changes, making the tuning of its overall loss threshold problematic – smaller values result in excessive clusters, while larger values may fail to detect group-specific concept drift altogether.
VI-D Effect of Window Size on Fairness and Performance
@@
Finally, we explore the influence of window size on both fairness and performance metrics in our FL framework. We examine how varying window sizes affect the system’s ability to adapt to concept drift while considering the trade-offs between overhead and model effectiveness. To evaluate the impact of window size, we conducted experiments across different window size configurations, ranging from retaining the entire history of client identities and local datasets to utilizing only a sliding window of the latest timesteps.
Figure 5 illustrates the results of this experiment. It can be observed that varying the window size had minimal effect on both fairness and performance across different configurations. Despite the reduction in historical data retention, the system maintained comparable levels of fairness and performance across varying window sizes.
Given the negligible impact of window size on fairness and performance, it is possible to mitigate the overhead associated with preserving the entire history of client identities and local datasets - clients can opt to maintain only a sliding window of the latest timesteps. By adopting this approach, clients can significantly reduce the computational and storage burden without compromising fairness or performance. Moreover, in certain scenarios, forgetting older data might actively aid in adaptating to concept drift, enhancing the system’s resilience to changing data distributions over time.
@@
Figure 5: Effect of window size on fairness and performance on the MNIST-GDrift dataset considering $\alpha=0.1$.
VII Conclusions and Future Work
@@
This work presents a pioneering effort in introducing and formalizing the problem of group-specific distributed concept drift within FL frameworks. Our experiments underscore the detrimental impact of this phenomenon on fairness, emphasizing the need for effective mechanisms to address evolving data distributions while upholding fairness principles. Our contributions include the formalization of group-specific concept drift and its distributed counterpart, an experimental framework for studying it, and the application of model clustering and fairness monitoring techniques to mitigate the challenges posed by this issue.
@@
We demonstrate that in the absence of group-specific distributed concept drift handling, fairness tends to decrease over time. In addition, for scenarios with high group imbalance, our findings indicate that overall loss and accuracy do not change significantly. However, there is a noticeable decrease in fairness due to the changing loss of the unprivileged group. In contrast, as the representation of the unprivileged group increases, both overall loss and the loss of the unprivileged group exhibit more noticeable changes. Our proposed algorithm, FairFedDrift, emerged as a practical solution for achieving fairness in FL amid group-specific distributed concept drift. By continuously monitoring group loss and adapting to changes in data distributions, FairFedDrift demonstrated its effectiveness in mitigating unfairness.
As we look to the future, several directions for further research emerge. Automating the hyperparameter selection process for FairFedDrift could streamline its application across diverse scenarios. Additionally, developing specialized fairness metrics tailored for evaluating concept drift in distributed learning settings would provide more nuanced insights into the evolving landscape of fairness.
@@
While our research aims to achieve fairness, it is essential to recognize that this objective may not be guaranteed in all use cases. The fairness metrics selected for evaluation and the choice of sensitive attributes are often contingent upon the application domain. Moreover, the binarization process employed to represent different groups of sensitive attributes may not capture the diversity within each group. These considerations underscore the complexity of fairness in machine learning and highlight the importance of ongoing refinement to address evolving ethical concerns.

### v4

@@ VI-A Results Across Multiple Datasets and Scenarios
VI-B Unveiling the Influence of $\alpha$ on Loss
We explore how varying the $\alpha$ parameter, which determines the size of the unprivileged group relative to the privileged group, impacts both the overall loss $\ell$ and group-specific loss $\ell_{S=0}$. We use FedAvg as the base algorithm to observe the behavior of the model under two different $\alpha$ values: 0.1 and 0.5. These values represent different levels of heterogeneity in group distributions, with lower $\alpha$ values indicating more heterogeneous groups and higher $\alpha$ values representing more homogeneous ones.
@@ VI-B Unveiling the Influence of $\alpha$ on Loss
Figure 3: Average clients’ loss ($\ell$) and group-specific ($S=0$) loss ($\ell_{S=0}$) results for the MNIST-GDrift dataset considering Scenario 4.1, $\alpha=0.1$ and $\alpha=0.5$.
@@
Figure 4: Clustering of FedDrift, FairFedDrift and Oracle on the MNIST-GDrift dataset, considering $\alpha=0.1$ and Scenario 4.1. Each cell represents the model ID at each client and timestep, while background colors denote the ground-truth concept.
Figure 3 shows the results of these experiments. For $\alpha=0.1$, the overall loss $\ell$ remains relatively stable over time, fluctuating only slightly. However, when we isolate the loss for the unprivileged group ($S=0$), we observe much more substantial variation. This indicates that even though the model’s overall performance does not appear to degrade, certain subsets of the population (in this case, the unprivileged group) are experiencing higher loss, which raises fairness concerns. This scenario highlights a significant challenge: relying on overall performance as the sole indicator of drift can obscure issues faced by smaller or less represented groups.
@@
FedDrift, designed to detect concept drift by monitoring loss changes, faces challenges in these situations. Since the overall loss remains stable, the algorithm may struggle to detect drift in specific subgroups such as the unprivileged group. Consequently, tuning the threshold $\delta$ in FedDrift becomes difficult - smaller values can lead to too many clusters, while larger values may fail to detect group-specific concept drift. This illustrates the limitations of FedDrift in addressing subtle changes affecting only a portion of the population, as previously demonstrated.
Conversely, for $\alpha=0.5$, both the overall loss and the unprivileged group’s loss exhibit more pronounced variations. In this case, the unprivileged group is better represented among the clients, leading to more noticeable loss changes. The fluctuations in both the overall model loss and the group-specific loss occur together, making it easier for FedDrift to detect drift and adjust accordingly. This scenario reflects a situation where drift detection is more effective, as group-specific loss changes significantly impact the overall performance.
The key takeaway from these experiments is that in scenarios with highly heterogeneous distributions, where the unprivileged group is less represented, concept drift can be difficult to detect using traditional drift detection methods that do not account for group-specific variations. In real-world applications, heterogeneous distributions are quite common, and unprivileged groups are typically in the minority. As such, detecting drift in these groups is important for maintaining fairness across the population. Models that rely solely on overall performance metrics risk overlooking fairness issues and disproportionately affecting underrepresented groups. This underscores the importance of designing drift detection algorithms, such as the one present in the FairFedDrift algorithm, that are sensitive to group-specific performance changes, even when overall metrics appear stable.
Figure 5: Model merge process over time in FairFedDrift on the MNIST-GDrift dataset, considering $\alpha=0.1$ and Scenario 4.1. Each node represents a global model at a specific timestep, with background colors denoting the original concepts in which the models where created. Edges indicate model merges based on group-specific loss thresholds. This Figure complements Figure 4, providing an aggregated view of how model clusters evolve and merge over time.
Figure 6: Effect of $\delta_{s}$ on fairness and performance across the MNIST-GDrift, FEMNIST-GDrift, Adult-GDrift and CIFAR-10-GDrift datasets evaluated under Scenario 4.1.
@@
Figure 4 presents the clustering outcomes achieved by FedDrift, FairFedDrift and Oracle on the MNIST-GDrift dataset considering $\alpha=0.1$ and Scenario 4.1. Although both FedDrift and FairFedDrift can detect concept drift in the third timestep, it can be observed in the fourth timestep that clients in the green concept in FedDrift prefer to associate themselves with the original model (‘0’), while FairFedDrift’s clients in the green concept associate themselves with models created in the previous timestep. While FairFedDrift’s approach to associating clients according to the sum of losses (line 8 in Algorithm 1) is important in this scenario, FedDrift lacks this mechanism, resulting in clients preferring the original model.
Furthermore, FairFedDrift effectively merges models corresponding to the same underlying concept (e.g., ‘1-3-4-5’). This behavior is further illustrated in Figure 5, which visualizes the model merge process over time. While Figure 4 highlights per-client model associations across timesteps, Figure 5 offers an additional view of how FairFedDrift merges models. This visual analysis demonstrates FairFedDrift’s ability to achieve more accurate and adaptive clustering, thereby offering a robust solution to the problem of group-specific distributed concept drift.
@@
Finally, FairFedDrift offers easier tuning in situations of group-specific concept drift compared to FedDrift. The loss of the unprivileged group undergoes significant variations over time, while the overall loss remains relatively stable. This makes tuning FedDrift more challenging, as it may struggle to detect significant loss changes, making the tuning of its overall loss threshold problematic - smaller values result in excessive clusters, while larger values may fail to detect group-specific concept drift altogether.
VI-D Effect of $\delta_{s}$ on Fairness and Performance
We explore the effect of tuning the hyperparameter $\delta_{s}$, which controls the acceptable loss difference for a group $s$ between two consecutive timestamps on a global model. $\delta_{s}$ also acts as a threshold for merging similar models, as described in Algorithm 1. The results of different values of $\delta_{s}$ are presented in Figure 6.
For the Adult-GDrift dataset, smaller values of $\delta_{s}$ (between 0.05 and 0.5) work best. In this case, concept drift occurs only through changes in specific smaller subgroups, leading to smoother variations in the loss. In contrast to the other datasets, where drift involves more noticeable label swaps, the drift in Adult-GDrift is more subtle. As a result, smaller $\delta_{s}$ thresholds are better suited to detect and respond to these drifts. On the other hand, for the image datasets (MNIST-GDrift, FEMNIST-GDrift, and CIFAR-10-GDrift), larger thresholds (around 1.25 to 1.75) are more appropriate. In this case, a higher $\delta_{s}$ prevents over-sensitivity to minor loss variations and stabilizes clustering.
Finally, with respect to fairness-utility trade-offs, it can be seen that when $\delta_{s}$ is very small, FairFedDrift may create too many clusters (even one per client) which reduces both fairness and accuracy due to insufficient data per model. As $\delta_{s}$ increases, clustering becomes more meaningful, leading to improvements in both metrics. However, if $\delta_{s}$ is set too high (e.g., MNIST-GDrift with $\alpha=0.05$ and $\delta_{s}=2.0$), drift may remain undetected, causing all clients to converge to a single model - effectively reducing the method to standard FedAvg. In such cases, utility may remain high due to larger training data, but fairness deteriorates, particularly when the unprivileged group is underrepresented.
VI-E Effect of Window Size on Fairness and Performance
@@
Finally, we introduce and examine the effect of window size on the performance of our framework. As explained before, the window size parameter determines how much past information is used during model assignment and concept drift detection. To evaluate the impact of window size, we conducted experiments across different window size configurations, ranging from retaining the entire history of client identities and local datasets to utilizing only a sliding window of the latest timesteps.
Figure 7 illustrates the results of this experiment. It can be observed that in this case maintaining a window size of 3 was sufficient, as increasing the window size beyond this had minimal impact on both fairness and performance across different configurations. Even with a reduced history of data retention, the system consistently achieved comparable levels of fairness and accuracy, demonstrating that a small window size strikes a good balance between computational efficiency and effective drift detection without sacrificing performance.
@@
Figure 7: Effect of window size on fairness and performance on the MNIST-GDrift dataset considering $\alpha=0.1$ and Scenario 4.1.
Given the negligible impact of window size on fairness and performance, it is possible to mitigate the overhead associated with preserving the entire history of client identities and local datasets - clients can opt to maintain only a sliding window of the latest timesteps. By adopting this approach, clients can significantly reduce the computational and storage burden without compromising fairness or performance. Moreover, in certain scenarios, forgetting older data might actively aid in adapting to concept drift, enhancing the system’s resilience to changing data distributions over time.
VI-F Fairness and Performance Across Varying Scenario Complexity
We analyze the results across two different scenarios (4.1 and 4.5 in Figure 2), which vary in complexity and drift patterns. Scenario 4.1 represents a simpler case with only three distinct concepts, while Scenario 4.5 presents a more challenging environment, featuring five distinct concepts that emerge in a less predictable, more random fashion. Some concepts even appear simultaneously for the first time, adding further complexity to the drift detection and adaptation.
Scenario Metric Algorithm
FedAvg FedDrift FairFedDrift Oracle
4.1 AEQ 0.84+-0.07 0.85+-0.09 0.91+-0.08 0.94+-0.05
OEQ 0.81+-0.06 0.82+-0.08 0.87+-0.08 0.89+-0.05
OPP 0.81+-0.06 0.82+-0.08 0.88+-0.08 0.90+-0.05
ACC 0.94+-0.01 0.96+-0.02 0.94+-0.04 0.95+-0.02
4.5 AEQ 0.80+-0.10 0.80+-0.09 0.90+-0.09 0.92+-0.05
OEQ 0.77+-0.09 0.78+-0.07 0.86+-0.09 0.88+-0.05
OPP 0.77+-0.10 0.77+-0.07 0.86+-0.09 0.89+-0.05
ACC 0.94+-0.02 0.95+-0.02 0.93+-0.04 0.94+-0.03
TABLE II: Results for two distinct scenarios presented in Figure 2 on the MNIST-GDrift dataset and $\alpha=0.1$.
The results, summarized in Table II, illustrate that while Scenario 4.1 was the easiest to handle due to its predictable concept changes, Scenario 4.5 was the most difficult due to the unpredictability and overlap of drifts. Despite this complexity, FairFedDrift consistently performed better compared to FedAvg and FedDrift in terms of fairness across both scenarios, maintaining a competitive accuracy level. Notably, even in the most challenging setting of Scenario 4.5, FairFedDrift maintained a high level of fairness. Moreover, FairFedDrift’s fairness and performance in both scenarios closely approached that of the idealized Oracle algorithm.
These results underscore the effectiveness of the proposed method in handling both simple and complex drift scenarios, demonstrating FairFedDrift’s capability to preserve fairness without significant accuracy loss, even under challenging conditions.
VII Discussion and Future Work
This paper introduces and formalizes the concept of group-specific distributed concept drift within FL, providing a framework to address fairness challenges in scenarios with evolving data distributions. While significant progress has been made in advancing fairness-aware FL, several directions remain open for further exploration, particularly in improving the scalability, efficiency, and practical deployment of FairFedDrift.
Automatic Hyperparameter Tuning
One promising area for future work is the automation of the hyperparameter selection process for group-specific thresholds. In its current form, these hyperparameters require manual tuning, which may limit the generalizability of the algorithm to different datasets and application scenarios. Developing adaptive methods for these parameters would enhance the flexibility and robustness of FairFedDrift, reducing the need for expert intervention and allowing for broader applicability across diverse use cases.
Fairness Metrics and Evaluation under Concept Drift
Existing fairness metrics do not fully account for the subtleties of concept drift in FL settings. There is a need to develop specialized metrics that can capture the evolving fairness dynamics as data distributions drift over time. These dynamic metrics should track fairness over time, providing insights into how well fairness-aware algorithms such as FairFedDrift can adapt to evolving conditions and ensure equitable outcomes across different groups.
Scalability and Communication Overhead
A key challenge for the real-world application of FairFedDrift is ensuring scalability and minimizing communication overhead. Monitoring group-specific losses and managing multiple global models can introduce significant computational and communication costs, particularly in large-scale federated systems. To address this, one potential approach is to use gradient similarity techniques to identify models with similar learning patterns and parameters. Additionally, developing novel mechanisms for dynamically selecting the best global model for each client based on their data, such as a cluster assignment model, could further enhance efficiency. Future research should also focus on analyzing various metrics to evaluate latency, communication efficiency, and computational overhead, considering the time, data exchanged, and computational resources required for each operation described in Section IV.
Real-World Federated Learning Datasets with Concept Drift
To further validate the robustness and applicability of FairFedDrift, future research should focus on real-world datasets that embody spacial and temporal concept drift. Developing and using datasets from sensitive domains that capture the complexities of group-specific concept drift in decentralized environments would be important for a realistic evaluation of FairFedDrift’s effectiveness in addressing fairness challenges in real-world FL settings.
VIII Conclusions
@@
This work presents a pioneering effort in introducing and formalizing the problem of group-specific concept drift within FL frameworks. Our experiments underscore the detrimental impact of this phenomenon on fairness, emphasizing the need for effective mechanisms to address evolving data distributions while upholding fairness principles. Our contributions include the formalization of group-specific concept drift and its distributed counterpart, an experimental framework for studying it, and the application of model clustering and fairness drift monitoring techniques to mitigate the challenges posed by this issue.
@@
We demonstrate that in the absence of group-specific distributed concept drift handling, fairness decreases. In addition, for scenarios with high group imbalance, our findings indicate that overall loss and accuracy do not change significantly. However, there is a noticeable decrease in fairness due to the changing loss of the unprivileged group. In contrast, as the representation of the unprivileged group increases, both overall loss and the loss of the unprivileged group exhibit more noticeable changes. Our proposed algorithm, FairFedDrift, emerged as a practical solution for achieving fairness in FL amid group-specific distributed concept drift. By continuously monitoring group-specific loss and adapting to changes in data distributions, FairFedDrift demonstrated its effectiveness in mitigating unfairness.
@@
While our research aims to achieve fairness, it is essential to recognize that this objective may not be guaranteed in all use cases. The fairness metrics selected for evaluation and the choice of sensitive attributes are often contingent upon the application domain. Moreover, the binarization process employed to represent different groups of sensitive attributes may not capture the diversity within each group. These considerations underscore the complexity of fairness in machine learning and highlight the importance of continual improvement to address evolving ethical concerns.